# EvRT-DETR: Latent Space Adaptation of Image Detectors for Event-based Vision

### v2

@@ Abstract
Event-based cameras (EBCs) have emerged as a bio-inspired alternative to traditional cameras, offering advantages in power efficiency, temporal resolution, and high dynamic range. However, the development of image analysis methods for EBCs is challenging due to the sparse and asynchronous nature of the data. This work addresses the problem of object detection for EBC cameras. The current approaches to EBC object detection focus on constructing complex data representations and rely on specialized architectures. We introduce I2EvDet (Image-to-Event Detection), a novel adaptation framework that bridges mainstream object detection with temporal event data processing. First, we demonstrate that a Real-Time DEtection TRansformer, or RT-DETR, a state-of-the-art natural image detector, trained on a simple image-like representation of the EBC data achieves performance comparable to specialized EBC methods. Next, as part of our framework, we develop an efficient adaptation technique that transforms image-based detectors into event-based detection models by modifying their frozen latent representation space through minimal architectural additions. The resulting EvRT-DETR model reaches state-of-the-art performance on the standard benchmark datasets Gen1 (mAP $+2.3$) and 1Mpx/Gen4 (mAP $+1.4$). These results demonstrate a fundamentally new approach to EBC object detection through principled adaptation of mainstream architectures, offering an efficient alternative with potential applications to other temporal visual domains. The code is available at: https://github.com/realtime-intelligence/evrt-detr.
@@ 1 Introduction
[…]
@@
Figure 1: Object Detection Performance vs Inference Time. A summary of object detection performance (COCO mAP) vs inference time (ms) of various models on 1Mpx automotive dataset. Circle size is proportional to the number of model parameters. Inference times are reported for NVIDIA T4 GPU. Our EvRT-DETR models achieve state-of-the-art accuracy while maintaining competitive inference speeds.
@@
Figure 2: Event-Based Camera Output Visualization. Sample frames from the 1Mpx dataset showing event polarity: red indicates positive events and blue indicates negative events.
@@
These unique characteristics of EBC data – sparsity, asynchronicity, and temporal nature – create fundamental challenges when applying traditional computer vision techniques. Early attempts to directly apply conventional image-based object detectors to simple event representations yielded poor performance [26], highlighting the gap between these domains. This led to two predominant research directions: (1) constructing sophisticated event representations that transform EBC data into formats compatible with existing detectors (e.g., [39]), or (2) designing specialized architectures tailored to handle the temporal nature of event data (e.g., [24]). A common assumption in these approaches is that EBC data processing requires either specialized representations or architectures due to the fundamental differences between event-based and frame-based vision, leading to development paths that often diverge from mainstream computer vision research.
@@
Meanwhile, mainstream object detection has seen remarkable progress, evolving from CNN-based architectures like YOLO [28] and Faster R-CNN [29] to transformer-based models like DETR [2]. Recent advances have culminated in highly efficient and accurate models such as the Real-Time DEtection TRansformer (RT-DETR) [36], which achieves exceptional performance while maintaining real-time inference speeds. Though primarily optimized for processing individual images, these state-of-the-art detectors offer powerful feature extraction capabilities that could potentially be leveraged for temporal data.
@@
This evolution in mainstream detection creates an opportunity to bridge the gap between conventional and event-based vision. Rather than assuming fundamental incompatibility between these domains, we explore a complementary direction: adapting state-of-the-art frame-based detectors to effectively process event data with minimal modifications. In this work, we investigate how RT-DETR can be strategically adapted to the event-based domain, demonstrating that modern object detection architectures can successfully work on event data and opening a new research direction.
[…]
@@
Building on this foundation, we introduce I2EvDet (Image-to-Event Detection), a technique that enables straightforward transformation of image-based object detectors into video-capable models. Our approach leverages latent space adaptation principles [16, 17] where we freeze the pre-trained RT-DETR model and insert a lightweight recurrent neural network (RNN) module within the encoder’s latent representation space. This design enables the model to capture temporal dynamics across frames while preserving the powerful representation capabilities of the original detector and adding minimal computational overhead. The resulting adaptation effectively leverages information across multiple time windows without requiring a specialized video detection architecture.
@@
The combined approach, which we call EvRT-DETR, achieves state-of-the-art performance on standard EBC benchmarks, outperforming specialized methods on both Gen1 [6] (mAP $+2.3$) and 1Mpx [26] (mAP $+1.4$) datasets. Importantly, these results are achieved using only standard modules from natural image and video analysis, without requiring domain-specific modifications. This demonstrates not only the effectiveness of our approach for EBC data but also suggests a broader paradigm for adapting image detectors to temporal domains. As shown in Figure 1, our model achieves superior accuracy while maintaining competitive inference speeds, making it practical for real-world applications.
@@
  • We show that mainstream object detectors like RT-DETR, when trained on simple image-like representations of EBC data using standard procedures, can achieve performance comparable to specialized EBC methods, offering a complementary approach to domain-specific architectures.
  • We propose I2EvDet, an efficient adaptation technique that transforms image-based detectors into video-capable models through minimal architectural additions to the frozen pre-trained model, with potential applications to various temporal visual domains.
  • We validate our approach in the challenging domain of event-based cameras, where our EvRT-DETR achieves state-of-the-art performance on the standard Gen1 (mAP $+2.3$) and 1Mpx (mAP $+1.4$) EBC benchmarks while maintaining competitive inference speeds.
@@ 2 Related Work
Research at the intersection of event-based cameras and object detection spans multiple domains in computer vision. Current approaches primarily focus on two directions: designing efficient representations for the unique properties of event data, and developing specialized neural architectures to process these representations. However, a third direction - adapting mainstream computer vision models to event data - remains underexplored despite its success in other domains. This section reviews relevant work across these areas, highlighting how our approach bridges the gap between specialized event-based methods and mainstream object detection advances.
@@ 2.1 EBC Data Representations
An event camera produces a stream of event data in the form $(t,p,x,y)$, where $t$ is the timestamp of an event, $p$ is the polarity (positive or negative) of the brightness change, and $(x,y)$ is the location of the event pixel. To simplify the analysis, the event data often are transformed into alternative representations, more suitable for conventional image analysis algorithms.
The simplest image-like representations of event camera data are Event Frames or 2D Histograms [8], where events within fixed time windows are accumulated into 2D frames of shape $(2,H,W)$ with polarity as the first dimension and spatial coordinates as the remaining dimensions. A natural evolution is the Stacked 2D Histogram, which further partitions each frame into $T$ time intervals, creating a representation of shape $(2,T,H,W)$ that is typically reshaped to $(2T,H,W)$ and treated as a natural image. While these representations are straightforward, directly applying existing computer vision algorithms to them yielded poor performance in the past [26].
To achieve better performance, other representations are being explored. For example, a Time Surface (TS) representation [39, 8] is an image-like representation where pixel values encode the time elapsed since the last event in a given pixel. Thus, unlike the simplest fixed time window representations, TS potentially can encode arbitrarily distant past. Empirically, it has been shown that using the TS representations offers significantly better performance on the object detection tasks compared to the 2D Histogram representations [26].
Many other EBC data representations also have been proposed. The recent ERGO-12 [39] work has developed efficient image-like representations, achieving state-of-the-art object detection performance on the Gen1 dataset. Meanwhile, other approaches reformulate object detection as a three-dimensional (3D) problem by treating events as points in spatial-temporal space or discretizing them into voxel grids [8]. While computationally heavier, these representations can preserve exact event timing, potentially enabling better handling of overlapping objects and complex motions.
@@ 2.2 EBC Object Detection Architectures
The most straightforward approach applies conventional image detection methods directly to 2D Histogram frames, but this yields relatively poor performance [26]. More advanced methods like Recurrent Vision Transformer (RVT) [11] and State Space Vision Transformer (S5-ViT-B) [40] extend this approach by incorporating temporal memory to capture sequential information, significantly improving detection quality compared to frame-by-frame processing.
The most recent family of methods attempts to use more efficient event representations to further improve the object detection performance. For example, 12-channel Event Representation through Gromov-Wasserstein Optimization (ERGO-12) [39] constructs an optimized image-like event representation and shows that the standard object detection methods perform exceptionally well on it. Alternatively, Asynchronous Spatio-Temporal Memory Network (ASTMNet) [19] and Adaptive Event Conversion (AEC) [23] attempt to combine both better event representation and use novel neural architectures architecture to expand state-of-the-art results.
Multiple other methods have been developed either treating EBC object detection as a 3D detection problem [30, 26, 10], exploiting benefits of neuromorphic architectures [3], or exploring various hybrid approaches [24, 25].
Figure 3: Overview of the I2EvDet framework applied to RT-DETR. Our approach adapts a pre-trained RT-DETR model to process temporal event data through minimal architectural modifications. The backbone and encoder components remain frozen during temporal adaptation, while trainable temporal modules are strategically inserted at multiple feature scales. Left: Event camera frames are processed through the frozen RT-DETR backbone and encoder. Center: Our temporal adaptation modules operate on multiscale features through residual connections (red pathways), enabling information flow across time steps while preserving spatial representation integrity. Right: Feature fusion and decoder stages produce final detections. The detailed view of the Temporal Adaptation Module (bottom right) shows the projection operations ($W_{i}^{down},W_{i}^{up}$) and recurrent memory module ($\mathbf{R}_{i}$) that enable efficient temporal modeling with minimal additional parameters.
@@ 2.3 Transformer-based Object Detection
Object detection has been revolutionized by transformer architectures in recent years. DEtection TRansformer (DETR) [2] was the first widely successful Transformer-based [33] model for object detection, offering a simple and elegant alternative to traditional CNN-based architectures [12, 28].
@@
Despite its innovative approach, the original DETR suffered from limitations including training instability [34], poor performance on small objects [2], and high computational costs during inference [38]. Subsequent research has addressed these issues through architectural improvements such as Feature Pyramid Networks (FPN), specialized decoder tokens, and deformable attention mechanisms [38, 2].
These incremental advances culminated in the Real-Time DEtection TRansformer (RT-DETR) [36], which achieves exceptional performance across all object scales with stable training and significantly faster inference times. RT-DETR represents the current state-of-the-art in efficient transformer-based detection and serves as an ideal foundation for adaptation to specialized domains such as event-based vision.
@@ 2.4 Model Adaptation Techniques
While specialized architectures dominate event-based vision, adapting mainstream architectures offers compelling advantages. Parameter-efficient methods [16, 14, 17] modify pre-trained models with minimal changes by adding compact modules to frozen backbones. These approaches have proven effective across language [7], vision [37], and multimodal systems [13], but remain underexplored for event-based vision despite their efficiency and transfer learning benefits.
Given the success of adaptation techniques in other domains and the power of transformer-based detection models, there is a clear opportunity to bridge these approaches for event-based vision. Our work explores this intersection, demonstrating that mainstream object detectors can be effectively adapted to EBC data through minimal modifications rather than requiring specialized architectures.
@@ 3 Methods
This section presents I2EvDet (Image-to-Event Detection), our adaptation framework for event-based object detection. Unlike approaches that rely on specialized architectures or complex representations, I2EvDet bridges mainstream object detection with temporal event data through strategic, minimal modifications. Our framework demonstrates how image-based detectors can be efficiently transformed into video-capable models while preserving their core strengths. We first describe our image-like representation of event data and then introduce the I2EvDet framework. Finally, we detail the technical implementation and discuss how our approach addresses the unique temporal dependencies in event camera data.
@@ 3.1 Event Representation
Our adaptation framework begins with transforming EBC data into an image-like representation that’s compatible with mainstream object detection models. We employ the Stacked 2D Histogram representation [8] as it provides a straightforward bridge between event data and conventional image processing architectures.
[…]
To construct a frame corresponding to an interval $[t_{0},t_{0}+T_{\text{frame}}]$, we create an intermediate stacked histogram $S$ from a set of events $\mathcal{E}$ in that interval:
@@
where $t_{i}=t_{0}+i\cdot T_{\text{bin}}$, $i\in[0,10)$ is the bin index, $(x,y)$ are spatial indices of the event in the EBC matrix, and $\mathbf{1}$ is an indicator function. Once a stacked histogram $S$ of shape $(2,10,H,W)$ is constructed, we merge the polarity and time dimensions to obtain an image-like 2D frame $F$ of shape $(20,H,W)$.

### v3

@@ Abstract
Event-based cameras (EBCs) have emerged as a bio-inspired alternative to traditional cameras, offering advantages in power efficiency, temporal resolution, and high dynamic range. However, development of image analysis methods for EBCs is challenging due to the sparse and asynchronous nature of the data. This work addresses the problem of object detection for EBC cameras. The current approaches to EBC object detection focus on constructing complex data representations and rely on specialized architectures. We introduce I2EvDet (Image-to-Event Detection), a novel adaptation framework that bridges mainstream object detection with temporal event data processing. First, we demonstrate that a Real-Time DEtection TRansformer, or RT-DETR, a state-of-the-art natural image detector, trained on a simple image-like representation of the EBC data achieves performance comparable to specialized EBC methods. Next, as part of our framework, we develop an efficient adaptation technique that transforms image-based detectors into event-based detection models by modifying their frozen latent representation space via minimal architectural additions. The resulting EvRT-DETR model reaches state-of-the-art performance on the standard benchmark datasets Gen1 (mAP +2.3+2.3) and 1Mpx/Gen4 (mAP +1.4+1.4). These results demonstrate a fundamentally new approach to EBC object detection through principled adaptation of mainstream architectures, offering an efficient alternative with potential applications to other temporal visual domains. The code is available at: https://github.com/realtime-intelligence/evrt-detr.
@@ 1 Introduction
Event-based cameras (EBCs) present a biologically inspired alternative to traditional frame-based cameras. Unlike conventional cameras that capture frames at predetermined intervals, EBC pixels operate asynchronously, generating data only when brightness changes exceed a threshold. This results in sparse, asynchronous data streams where each event is a tuple containing pixel location, timestamp, and polarity (whether brightness increased or decreased), as illustrated in Figure 2. This approach offers remarkable advantages: low power consumption (as low as $\mathrm{1}\mathrm{0}$ mW), reduced data transfer rates, exceptional temporal resolution (on the order of $\mu s$), and high dynamic range ($\mathrm{1}\mathrm{0}\mathrm{0}$ dB). These properties have led to widespread adoption in autonomous driving, robotics, and wearable electronics [8].
@@
Figure 1: Object Detection Performance vs. Inference Time. A summary of object detection performance ($CO$ mAP) versus inference time (ms) of various models on the 1Mpx automotive dataset. Circle size is proportional to the number of model parameters. Inference times are reported for NVIDIA T4 GPU. Our EvRT-DETR models achieve state-of-the-art accuracy while maintaining competitive inference speeds.
@@
Figure 2: Event-based Camera Output Visualization. Sample frames from the 1Mpx dataset showing event polarity. Red indicates positive events, and blue denotes negative events.
@@
These unique characteristics of EBC data – sparsity, asynchronicity, and temporal nature – create fundamental challenges when applying traditional computer vision techniques. Early attempts to directly apply conventional image-based object detectors to simple event representations yielded poor performance [28], highlighting the gap between these domains. This led to two predominant research directions: (1) constructing sophisticated event representations that transform EBC data into formats compatible with existing detectors (e.g., [41]) or (2) designing specialized architectures tailored to handle the temporal nature of event data (e.g., [26]). A common assumption in these approaches is that EBC data processing requires either specialized representations or architectures due to the fundamental differences between event- and frame-based vision, leading to development paths that often diverge from mainstream computer vision research.
@@
Meanwhile, mainstream object detection has seen remarkable progress, evolving from convolutional neural network (CNN)-based architectures, e.g., YOLO [30] and Faster R-CNN [31], to transformer-based models like DETR (DEtection TRansformer) [2]. Recent advances have culminated in highly efficient and accurate models, such as the Real-Time DEtection TRansformer (RT-DETR) [38] that achieves exceptional performance while maintaining real-time inference speeds. Though primarily optimized for processing individual images, these state-of-the-art detectors offer powerful feature extraction capabilities that potentially could be leveraged for temporal data.
@@
This evolution in mainstream detection creates an opportunity to bridge the gap between conventional and event-based vision. Rather than assuming fundamental incompatibility between these domains, we explore a complementary direction: adapting state-of-the-art frame-based detectors to process event data effectively with minimal modifications. In this work, we investigate how RT-DETR can be strategically adapted to the event-based domain, demonstrating that modern object detection architectures can successfully work on event data and opening a new research direction.
Our approach proceeds in two stages to systematically adapt RT-DETR to event data. First, we train RT-DETR directly on simple image-like representations of EBC data using standard training procedures without special modifications. This straightforward approach yields surprising results – even when processing only fixed $50\text{\,}\mathrm{m}\mathrm{s}$ time frames, the detector achieves performance comparable to specialized EBC-specific methods explicitly designed for video object detection on EBC data. This is particularly unexpected as these specialized methods deliberately exploit the temporal nature of EBC data and process information from the distant past, while our approach performs detection using only a single fixed time window (frame) of $50\text{\,}\mathrm{m}\mathrm{s}$, similar to traditional frame-based cameras.
@@
Building on this foundation, we introduce I2EvDet (Image-to-Event Detection), a technique that enables straightforward transformation of image-based object detectors into video-capable models. This approach leverages latent space adaptation principles [17, 18] where we freeze the pre-trained RT-DETR model and insert a lightweight recurrent neural network (RNN) module within the encoder’s latent representation space. This design allows the model to capture temporal dynamics across frames while preserving the powerful representation capabilities of the original detector and adding minimal computational overhead. The resulting adaptation effectively leverages information across multiple time windows without requiring a specialized video detection architecture.
@@
The combined approach, called EvRT-DETR, achieves state-of-the-art performance on standard EBC benchmarks, outperforming specialized methods on both Gen1 [6] (mAP +2.3+2.3) and 1Mpx [28] (mAP +1.4+1.4) datasets. Importantly, the results are achieved using only standard modules from natural image and video analysis without requiring domain-specific modifications. This demonstrates not only the effectiveness of our approach for EBC data, but also suggests a broader paradigm for adapting image detectors to temporal domains. As shown in Figure 1, our model achieves superior accuracy while maintaining competitive inference speeds, making it practical for real-world applications.
@@
  • We show that when trained on simple image-like representations of EBC data using standard procedures, mainstream object detectors like RT-DETR can achieve performance comparable to specialized EBC methods, offering a complementary approach to domain-specific architectures.
  • We propose I2EvDet, an efficient adaptation technique that transforms image-based detectors into video-capable models through minimal architectural additions to the frozen pre-trained model and with potential applications to various temporal visual domains.
  • We validate our approach in the challenging domain of event-based cameras, where EvRT-DETR achieves state-of-the-art performance on the standard Gen1 (mAP +2.3+2.3) and 1Mpx (mAP +1.4+1.4) EBC benchmarks while maintaining competitive inference speeds.
@@ 2 Related Work
Research at the intersection of event-based cameras and object detection spans multiple domains in computer vision. Current approaches primarily focus on two directions: designing efficient representations for the unique properties of event data and developing specialized neural architectures to process these representations. However, despite its success in other domains, a third direction – adapting mainstream computer vision models to event data – remains underexplored. This section reviews relevant work across these areas, highlighting how our approach bridges the gap between specialized event-based methods and mainstream object detection advances.
@@ 2.1 EBC Data Representations
An event camera produces a stream of event data in the form $(t,p,x,y)$, where $t$ is the timestamp of an event, $p$ is the polarity (positive or negative) of the brightness change, and $(x,y)$ is the location of the event pixel. To simplify the analysis, the event data often are transformed into alternative representations more suitable for conventional image analysis algorithms.
The simplest image-like representations of event camera data are Event Frames or 2D Histograms [8], where events within fixed time windows are accumulated into two-dimensional frames of shape $(2,H,W)$ with polarity as the first dimension and spatial coordinates as the remaining dimensions. A natural evolution is the Stacked 2D Histogram, which further partitions each frame into $T$ time intervals, creating a representation of shape $(2,T,H,W)$ that typically is reshaped to $(2T,H,W)$ and treated as a natural image. While these representations are straightforward, directly applying existing computer vision algorithms to them has yielded poor performance in the past [28].
To improve performance, other representations are being explored. For example, a Time Surface (TS) representation [41, 8] is an image-like representation where pixel values encode the time elapsed since the last event in a given pixel. Thus, unlike the simplest fixed time window representations, TS potentially can encode arbitrarily distant past. Empirically, it has been shown that using the TS representations offers significantly better performance on the object detection tasks compared to the 2D Histogram representations [28].
Many other EBC data representations also have been proposed. The recent ERGO-12 (Event Representation through Gromov-Wasserstein Optimization) [41] work has developed efficient image-like representations, achieving state-of-the-art object detection performance on the Gen1 dataset. Meanwhile, other approaches reformulate object detection as a three-dimensional (3D) problem by treating events as points in spatial-temporal space or discretizing them into voxel grids [8]. While computationally heavier, these representations can preserve exact event timing, potentially enabling better handling of overlapping objects and complex motions.
@@ 2.2 EBC Object Detection Architectures
The most straightforward approach applies conventional image detection methods directly to 2D Histogram frames, but this yields relatively poor performance [28]. More advanced methods, such as Recurrent Vision Transformer (RVT) [12] and State Space Vision Transformer (S5-ViT-B) [42], extend this approach by incorporating temporal memory to capture sequential information, significantly improving detection quality compared to frame-by-frame processing.
The most recent family of methods attempts to use more efficient event representations to further improve object detection performance. For example, 12-channel ERGO-12 [41] constructs an optimized image-like event representation and shows that the standard object detection methods perform exceptionally well on it. Alternatively, Asynchronous Spatio-Temporal Memory Network (ASTMNet) [20] and Adaptive Event Conversion (AEC) [25] attempt to combine better event representation while using novel neural architectures to expand state-of-the-art results.
Other methods have been developed that treat EBC object detection as a 3D detection problem [32, 28, 11], exploit benefits of neuromorphic architectures [3], or explore various hybrid approaches [26, 27].
Figure 3: Overview of the I2EvDet Framework Applied to RT-DETR. This approach adapts a pre-trained RT-DETR model to process temporal event data using minimal architectural modifications. The backbone and encoder components remain frozen during temporal adaptation, while trainable temporal modules are strategically inserted at multiple feature scales. Left: Event camera frames are processed through the frozen RT-DETR backbone and encoder. Center: Our temporal adaptation modules operate on multiscale features through residual connections (red pathways), enabling information flow across time steps while preserving spatial representation integrity. Right: Feature fusion and decoder stages produce final detections. The detailed view of the Temporal Adaptation Module (bottom right) shows the projection operations ($W_{i}^{down},W_{i}^{up}$) and recurrent memory module ($\mathbf{R}_{i}$) that enable efficient temporal modeling with minimal additional parameters.
@@ 2.3 Transformer-based Object Detection
Object detection has been revolutionized by transformer architectures in recent years. DETR [2] was the first widely successful transformer-based [35] model for object detection, offering a simple and elegant alternative to traditional CNN-based architectures [13, 30].
@@
Despite its innovative approach, the original DETR suffered from limitations, including training instability [36], poor performance on small objects [2], and high computational costs during inference [40]. Subsequent research has addressed these issues through architectural improvements such as Feature Pyramid Networks (FPN), specialized decoder tokens, and deformable attention mechanisms [40, 2].
These incremental advances have culminated in the RT-DETR [38], which achieves exceptional performance across all object scales with stable training and significantly faster inference times. RT-DETR represents the current state-of-the-art in efficient transformer-based detection and serves as an ideal foundation for adaptation to specialized domains, such as event-based vision.
@@ 2.4 Model Adaptation Techniques
While specialized architectures dominate event-based vision, adapting mainstream architectures offers compelling advantages. Parameter-efficient methods [17, 15, 18] modify pre-trained models with minimal changes by adding compact modules to frozen backbones. These approaches have proven effective across language [7], vision [39], and multimodal systems [14], but they remain underexplored for event-based vision despite their efficiency and transfer learning benefits.
Given the success of adaptation techniques in other domains and the power of transformer-based detection models, there is a clear opportunity to bridge these approaches for event-based vision. Our work explores this intersection, demonstrating that mainstream object detectors can be effectively adapted to EBC data with minimal modifications rather than requiring specialized architectures.
@@ 3 Methods
This section presents I2EvDet, our adaptation framework for event-based object detection. Unlike approaches that rely on specialized architectures or complex representations, I2EvDet bridges mainstream object detection with temporal event data via strategic, minimal modifications. Our framework demonstrates how image-based detectors can be efficiently transformed into video-capable models while preserving their core strengths. We first describe our image-like representation of event data then introduce the I2EvDet framework. Finally, we detail the technical implementation and discuss how our approach addresses the unique temporal dependencies in event camera data.
@@ 3.1 Event Representation
Our adaptation framework begins with transforming EBC data into an image-like representation that is compatible with mainstream object detection models. We employ the Stacked 2D Histogram representation [8] as it provides a straightforward bridge between event data and conventional image processing architectures.
To ensure direct comparability with existing methods, we adopt the frame construction parameters from RVT [12], which allows us to isolate the impact of our architectural choices from data pre-processing. Specifically, we take the stream of events and partition it into a series of consecutive fixed time windows of $T_{\text{frame}}=$50\text{\,}\mathrm{ms}$\$. Each such window corresponds to a single frame. Next, we further subdivide each frame into 10 intervals of $T_{\text{bin}}=$5\text{\,}\mathrm{ms}$\$.
To construct a frame corresponding to an interval $t_{0},t_{0}+T_{\text{frame}}]$, we create an intermediate stacked histogram $S$ from a set of events $\mathcal{E}$ in that interval:
@@
where $e=(t_{e},p_{e},x_{e},y_{e})$ represents an individual event with time stamp $t_{e}$, polarity $p_{e}$, and spatial coordinates $(x_{e},y_{e})$; $t_{i}=t_{0}+i\cdot T_{\text{bin}}$, $i\in[0,10)$ is the histogram bin index; $(x,y)$ are the spatial indices in the output histogram; $\delta_{a}^{b}$ is the Kronecker delta function; and $\mathbf{1}$ is an indicator function. Once a stacked histogram $S$ of shape $(2,10,H,W)$ is constructed, we merge the polarity and time dimensions to obtain an image-like 2D frame $F$ of shape $(20,H,W)$.